OpenStudy (junyang96):

Probability A box contains 5 red bulbs, 4 blue bulbs and 3 yellow bulbs. 3 bulbs are selected randomly from the box. Find the probability that all three bulbs are of different colours, given that one of them is yellow

OpenStudy (anonymous):

the probability of a yellow is 3\12 so most likely they wood be red

OpenStudy (junyang96):

@gorv

OpenStudy (junyang96):

@dan815

OpenStudy (anonymous):

If there are 12 bulbs and you select 3. The probability would be 3/12.

OpenStudy (junyang96):

No it's not that simple

OpenStudy (junyang96):

Answer is 5/9

OpenStudy (junyang96):

@satellite73

OpenStudy (anonymous):

hi we can do this

OpenStudy (junyang96):

I suppose it's a conditional probability. Isn't it?

OpenStudy (anonymous):

you started with 12 light bulbs, but you know one is yellow, so now it is like having 11 light bulbs and you have to pick 2 yes, it is conditional

OpenStudy (anonymous):

so change the problem to you have 5 red, 4 blue, 2 yellow you select 2 what is the probability that one is red and one is blue

OpenStudy (junyang96):

\[(5C1\times4C1)/11C2\]

OpenStudy (anonymous):

yeah that should do it easier to write \[\frac{5\times 4}{55}\]

OpenStudy (junyang96):

So is this the final answer?

OpenStudy (anonymous):

unless i miss my guess it is \(\frac{4}{11}\)

OpenStudy (junyang96):

49 junyang96 Medals 0 So here's my tutor's explanation. the Probability P( all 3 bulbs are of different colours| one yellow) =P(all 3 bulbs are of different colours bulbs AND one yellow)/P(one yellow) and since all 3 bulbs are of different colours is the subset of one yellow so =P(all 3 bulbs are of different colours)/P(one yellow) =[5C1*4C1*3C1/12C3]/[3C1*9C2/12C3] =5/9

OpenStudy (junyang96):

@ganeshie8

ganeshie8 (ganeshie8):

looks you have a solution ?

OpenStudy (junyang96):

I just want to know if which answer is correct. 5/9 or 4/11.

OpenStudy (junyang96):

satellite's explanation sounds reasonable too

ganeshie8 (ganeshie8):

your tutor is conditioning on : "exactly one bulb being yellow"

OpenStudy (junyang96):

yes

ganeshie8 (ganeshie8):

but that was not mentioned in the problem statement right ?

ganeshie8 (ganeshie8):

how do u know the condition is "exactly one bulb is yellow" or "atleast one bulb is yellow" ?

ganeshie8 (ganeshie8):

you get different answers based on how you interpret the problem

ganeshie8 (ganeshie8):

do you see how the phrasing is a bit ambiguous ?

ganeshie8 (ganeshie8):

your tutor is correct if we assume "exactly one bulb is yellow" satellite is correct if we assume "atleast one bulb is yellow"

OpenStudy (junyang96):

Thanks for the explanation. I suppose the question is saying exactly one bulb.

ganeshie8 (ganeshie8):

np :) lets fix the question then and work it again real quick : A box contains 5 red bulbs, 4 blue bulbs and 3 yellow bulbs. 3 bulbs are selected randomly from the box. Find the probability that all three bulbs are of different colours, given that `exactly one of them is yellow `

ganeshie8 (ganeshie8):

A : Exactly one of them is yellow B : All of them different

ganeshie8 (ganeshie8):

\[\rm P( B|A) = \dfrac{p(A \cap B)}{P(A)}\] you're using this formula, right ?

OpenStudy (junyang96):

yes

ganeshie8 (ganeshie8):

\[\rm P(A) = \dfrac{\binom{3}{1}\binom{9}{2}}{\binom{12}{3}}\] You can choose one yellow ball from the available 3 yellow balls in \(\binom{3}{1}\) ways The remaining two balls can be selected in \(\binom{9}{2}\) ways

ganeshie8 (ganeshie8):

\[\rm P(A\cap B) = \dfrac{\binom{5}{1}\binom{4}{1}\binom{3}{1}}{\binom{12}{3}}\] You're just picking one ball from each of the available colors

ganeshie8 (ganeshie8):

take the ratio

OpenStudy (junyang96):

thanks